p vector ( p vector + q vector + r vector ) = p sqare . find angle between p and q

Grade:11

1 Answers

Eshan
5 years ago
Dear student,

Since\vec{p}.(\vec{p}+\vec{q}+\vec{r})=p^2

we have,p^2+\vec{p}.(\vec{q}+\vec{r})=p^2\implies \vec{p}.(\vec{q}+\vec{r})=0

Therefore, all we can comment is that\vec{p}must be perpendicular to\vec{q}+\vec{r}.